If you are given the points (2,4), (x,8) with a slope = -2, how do you find the value of x?
Answer by stanbon(75887)

If you are given the points (2,4), (x,8) with a slope = -2, how do you find the value of x?
-----
slope = (8-4)/(x-2) = -2
---
4/(x-2) = -2
Cross-multiply:
-2(x-2) = 4
x-2 = -2
x = 0
===========
Cheers,
Stan H.